# Faster Quantum Simulation Of Markovian Open Quantum Systems Via Randomisation

### v2

@@ 4 Second Order Randomised Product Formula
The second order randomised SLT product formula is not much more complicated than its deterministic counterpart. It is constructed by considering a convex sum of all permutations of the exponentials in the second order SLT product formula. More precisely, consider the symmetric group $\mathrm{Sym}(M)$ which is the group of all permutations of the elements of the set $\{1,...,M\}$. For any permutation $\sigma\in\mathrm{Sym}(M)$ we define,
$S_{2}^{\sigma}(\tau):=\prod_{j=1}^{M}e^{\frac{\tau}{2}\hat{\mathcal{L}}_{\sigma(j)}}\prod_{j=M}^{1}e^{\frac{\tau}{2}\hat{\mathcal{L}}_{\sigma(k)}},$ (65)
which is a second order SLT product formula whose exponentials are permuted by the permutation $\sigma\in\mathrm{Sym}(M)$. Using this, we can construct the approximation to the total channel $T_{\tau}$ by taking a convex combination of $S_{2}^{\sigma}$ for all $\sigma\in\mathrm{Sym}(M)$,
@@
where the $1/M!$ is present because $|\mathrm{Sym}(M)|=M!$. The following theorem shows the error bound and the bound on $N$ for the second order randomised SLT product formula. The proof of this theorem is done in a similar way to the proof of the randomised product formulas in [11]. However, it also relies on the mixing lemma [28, 29], which as stated before is not applicable to open quantum systems, therefore we give a more direct proof.
@@
Given the generator $\mathcal{L}$ as in equation (10) of a quantum channel $T_{t}$ and some time $t\geq 0$. Define the second order randomised product formula as in equation (66). Let $N\in\mathbb{N}$ and $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$. Then,
$\left\|T_{t}-S_{2}^{(ran)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\epsilon,$ (67)
where,
$\epsilon\geq\frac{(2\Lambda t)^{3}M^{2}}{N^{2}},$ (68)
and,
$N\geq\frac{(2\Lambda t)^{3/2}M}{\epsilon^{1/2}}.$ (69)
In proving the following theorem we will need to Taylor expand the formula $S_{2}^{(ran)}$. However this may be a complicated and challenging task to do directly. Instead, we want to consider what an arbitrary order term looks like in the expansion. The following lemma, which we shall call the randomisation lemma, will tell us what the $s$-th order non-degenerate term looks like in the expansion of $S_{2}^{(ran)}$, where $0\leq s\leq M.$ We use the word non-degenerate to describe a product of constituent generators ^$\hat{\mathcal{L}}_{k}$ which is pairwise different. To make out calculations easier we rewrite the second order randomised formula $S_{2}^{(ran)}$ in the following general way
@@
The last sum in equation (4) is a permutation of all pairwise different $m_{1},...,m_{s}$ and we observe that for a fixed $m_{1},...,m_{s}$ there are $(M-s)!$ ways we can permute the rest of the indicies so that $m_{1},...,m_{s}$ is unchanged. Therefore, we remove this sum and add a factor $(M-s)!$, leading to the following expression for the $s$-th order non-degenerate term:
@@
The $s$-th order term in the Taylor expansion of $T_{\tau}$ is
$\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\end{subarray}}^{M}\hat{\mathcal{L}}_{m_{1}}...\hat{\mathcal{L}}_{m_{s}}$ (88)
and the non-degenerate part of (88) is given by Lemma 2. By bounding the diamond norm of the $s$-th order term and the non-degenerate part and taking their difference we can obtain a bound on the norm of the degenerate term. The bound on the norm of the $s$-th order term of $T_{\tau}$ is
$\left\|\frac{\tau^{s}}{s!}\sum_{m_{1},...,m_{s}=1}^{M}\hat{\mathcal{L}}_{m_{1}}...\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond}$ $\leq\frac{\tau^{s}}{s!}\sum_{m_{1},...,m_{s}=1}^{M}\left\|\hat{\mathcal{L}}_{m_{1}}\right\|_{\diamond}...\left\|\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond},$
$\leq\frac{\tau^{s}}{s!}\sum_{m_{1},...,m_{s}=1}^{M}\Lambda^{s},$
$=\frac{\tau^{s}}{s!}\Lambda^{s}M^{s}.$ (89)
Now we can bound the $s$-th order non-degenerate term of $T_{\tau}$ as
$\left\|\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{pairwise different}\end{subarray}}^{M}\hat{\mathcal{L}}_{m_{1}}...\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond}$ $\leq\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{pairwise different}\end{subarray}}^{M}\left\|\hat{\mathcal{L}}_{m_{1}}\right\|_{\diamond}...\left\|\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond},$
$\leq\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{pairwise different}\end{subarray}}^{M}\Lambda^{s},$
$=\frac{\tau^{s}}{s!}\Lambda^{s}[M(M-1)...(M-(s+1))],$ (90)
where the last equality is obtained by computing the sum which is equal to the number of all possible ways we can arrange $m_{1},...,m_{s}$ so that they are pairwise different. This implies that the bound on the norm of the degenerate term of $T_{\tau}$ is
$\frac{\tau^{s}\Lambda^{s}}{s!}[M^{s}-M(M-1)...(M-(s+1))].$ (91)
The norm of the $s$-th order degenerate term of $S_{2}^{(ran)}(\tau)$ is bounded in the same way as the norm of the $s$-th order degenerate term of $T_{\tau}$. We consider the $s$-th order non-degenerate term of $S_{2}^{(ran)}(\tau)$ in equation (72) calculated in Lemma 1. The bound on the norm is
$\left\|\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{pairwise different}\end{subarray}}^{M}\hat{\mathcal{L}}_{m_{1}}...\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond}\leq\frac{(\Lambda\tau)^{s}}{s!}M(M-1)...(M-(s+1)).$ (92)
To bound the $s$-th order term of $S_{2}^{(ran)}(\tau)$, we first, consider the general expression for $S_{2}^{(ran)}(\tau)$ in equation (4), replace each summand of the generator $\mathcal{L}$ by $\Lambda$ and replace each $q_{k}$ by 1/2. Next, we expand all exponentials into their Taylor series and extract the $s$-th order term. In other words, we extract the $s$-th order term of
$\frac{1}{M!}\sum_{\sigma\in\mathrm{Sym}(M)}\exp(M\Lambda\tau)$ (93)
to get,
$\frac{(M\Lambda\tau)^{s}}{s!}.$ (94)
Taking the difference of (94) and (92) yields the bound on the norm of the $s$-th order degenerate term of $S_{2}^{(ran)}(\tau)$. That is
$\frac{(\Lambda\tau)^{s}}{s!}[M^{s}-M(M-1)...(M-(s+1))].$ (95)
@@
Given a generator $\mathcal{L}$ as in equation (10) and the ideal evolution $T_{\tau}$ for some small time step $\tau\geq 0$ as well as the second order randomised formula $S_{2}^{(ran)}(\tau)$ defined in equation (66). The error of the approximation,
$T_{\tau}-S_{2}^{(ran)}(\tau)=\exp(\tau\mathcal{L})-\frac{1}{M!}\sum_{\sigma\in\mathrm{Sym}(M)}S_{2}^{\sigma}(\tau),$ (96)
is at most
$\begin{cases}0,\text{ for }0\leq s\leq 2,\\ \\ \frac{(\Lambda\tau)^{s}M^{s-1}}{(s-2)!},\text{ for }s>2.\end{cases}$ (97)
@@
Now we need to obtain a bound for the factor $[M^{s}-M(M-1)...(M-(s+1))]$. Using [11] we can obtain the following bound
@@ 5 The QDRIFT Channel For Simulating OQS
Given the generator $\mathcal{L}$ as in equation (8) of a quantum channel $T_{t}$ and some time $t\geq 0$. Let $N\in\mathbb{N}$ and $\Omega:=\max_{k}\left\|\mathcal{L}_{k}\right\|_{\diamond}$. Then,
$\left\|T_{t}-\left(\mathcal{E}^{(QD)}_{\omega}\right)^{N}\right\|_{\diamond}=\left\|T_{t}-\left(\sum_{k=1}^{M}p_{k}e^{\omega\mathcal{L}_{k}}\right)^{N}\right\|_{\diamond}\leq\epsilon$ (114)
@@
$\epsilon\geq\frac{t^{2}\Gamma^{2}\Omega^{2}}{N}$ (115)
@@
$N\geq\frac{t^{2}\Gamma^{2}\Omega^{2}}{\epsilon}.$ (116)
@@ 6.1.1 Implementation of $S_{1}^{(ran)}$ with CS
To implement the first order randomised product formula $S_{1}^{(ran)}(\tau)^{N}$, we need $N$ applications of $S_{1}^{(ran)}(\tau)$. However, the definition of $S_{1}^{(ran)}(\tau)$ in (30) tells us that when we apply it to a state $\rho(0)$ it will apply the channel $S_{1}^{\rightarrow}(\tau)$ with probability $1/2$ and $S_{1}^{\leftarrow}(\tau)$ with probability $1/2$. This gives us a way to construct a gate set for $S_{1}^{(ran)}(\tau)$. If we define some random variable $j_{l}\in\{0,1\}$ for $l=1,...N$ where $p(j_{l}=0)=1/2=p(j_{l}=1)$ then, by assigning $S_{1}^{\rightarrow}(\tau)\equiv S_{1}^{(0)}(\tau)$ and $S_{1}^{\leftarrow}(\tau)\equiv S_{1}^{(1)}(\tau)$, we can construct a gate set, denoted by $G_{1}^{(ran)}$, by iteratively sampling each $j_{l}$ and appending $S_{1}^{(j_{l})}(\tau)$ to $G_{1}^{(ran)}$. The gate set can then be applied to an initial state $\rho(0)$ to output the state $\tilde{\rho}(t)$ which approximates $\rho(t)$ to an precision $\epsilon=(Mt\Lambda)^{3}/3N^{2}$. The pseudo code for the algorithm that can construct this gate set is shown in Algorithm 1.
@@
Input: A list of terms from the generator $\mathcal{L}=\sum_{k=1}^{M}\hat{\mathcal{L}}_{k}$ i.e. {^$\{\hat{\mathcal{L}}_{1},...,\hat{\mathcal{L}}_{M}\}$. A classical oracle function called SAMPLE() that returns a value $j_{l}\in\{0,1\}$ from the distribution $p(j_{l}=0)=1/2$ and $p(j_{l}=1)=1/2$. A target precision $\epsilon$ and a simulation time $t\geq 0$.
@@
2:$N\leftarrow\left\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\ \right\rceil$
3:$\tau\leftarrow t/N$
4:$l\leftarrow 0$
5:$G_{1}^{(ran)}=\{\}$
@@
We want to obtain the gate complexity for the circuit implemented by $G_{1}^{(ran)}$. To do this, we count the number of simple channels, which in this case is the number of exponentials, $\exp(\tau\hat{\mathcal{L}}_{k})$, in the gate set. Since there are $\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ TS product formulas in $G_{1}^{(ran)}$, each containing $M$ exponentials, the gate complexity scales as $O(M^{5/2}(t\Lambda)^{3/2}/\sqrt{3\epsilon})$.
@@ 6.1.2 Implementation of $S_{2}^{(ran)}$ with CS
To the implement the second order randomised formula $S_{2}^{(ran)}(\tau)^{N}$, we need $N$ applications of $S_{2}^{(ran)}(\tau)$ as in (4). By definition $S_{2}^{(ran)}(\tau)$ applies $S_{2}^{\sigma}(\tau)$ for some permutation $\sigma\in\mathrm{Sym}(M)$ with a probability $1/M!$. But since the sum in (4) is over all possible permutations in $\mathrm{Sym}(M)$ it is equally likely that we apply a $S_{2}^{\sigma}(\tau)$ for any permutation $\sigma$. So to construct a gate set that will implement $S_{2}^{(ran)}(\tau)^{N}$ we define $\sigma_{l}$, for $l=1,2,...,N$, which will be some permutation from $\mathrm{Sym}(M)$. Then we define an oracle function called SAMPLE_PERMUTATION() which will sample from $\mathrm{Sym}(M)$, with every permutation in $\mathrm{Sym}(M)$ having a probability $1/M!$ of being sampled, and returning one of the permutations. We denote the gate set that we apply to initial state $\rho(0)$ as $G_{2}^{(ran)}$. Initially this set will be empty. Then, we iteratively obtain $\sigma_{l}$ using the function SAMPLE_PERMUTATION() i.e. $\sigma_{l}=$SAMPLE_PERMUTATION(), and append $S_{2}^{\sigma_{l}}$ to $G_{2}^{(ran)}$ repeating this process for $l=1,2,...,N$. Then, we can apply $G_{2}^{(ran)}$ to some initial sate $\rho(0)$ and the output will be the state $\tilde{\rho}(t)$ which approximates $\rho(t)$ to a precision $\epsilon=(\Lambda t)^{3}M^{2}/N^{2}$. Algorithm 2, shows the pseudo code for how one can $G_{2}^{(ran)}$. Figure 3, shows the quantum circuit for appliyng each gate from $G_{2}^{(ran)}$ where each $\sigma_{l}$ is some permutation sampled from $\mathrm{Sym}(M)$. To see how the gate complexity for a circuit constructed with $G_{2}^{(ran)}$ scales, we count the number of exponentials. We will have $\lceil(\Lambda t)^{3/2}M/\sqrt{\epsilon}\rceil$ TS product formulas in $G_{2}^{(ran)}$ and for each of these there are $2M$ exponentials. The gate complexity will then scale as $O((\Lambda t)^{3/2}M^{2}/\sqrt{\epsilon})$.
@@
Input: A list of terms from the generator $\mathcal{L}=\sum_{k=1}^{M}\hat{\mathcal{L}}_{k}$ i.e. {^$\{\hat{\mathcal{L}}_{1},...,\hat{\mathcal{L}}_{M}\}$. A classical oracle function called SAMPLE_PERMUTATION() that returns a permutation from $\mathrm{Sym}(M)$, with each permutation having the probability $1/M!$ of being sampled. A target precision $\epsilon$ and a simulation time $t\geq 0$.
@@
2:$N\leftarrow\left\lceil(\Lambda t)^{3/2}M/\sqrt{\epsilon}\ \right\rceil$
3:$\tau\leftarrow t/N$
4:$l\leftarrow 0$
5:$G_{2}^{(ran)}=\{\}$

### v3

@@ 4 Second Order Randomised Product Formula
The second order randomised Trotter-Suzuki product formula is not much more complicated than its deterministic counterpart. It is constructed by considering a convex sum of all permutations of the exponentials in the second order Trotter-Suzuki product formula. More precisely, consider the symmetric group $\mathrm{Sym}(M)$ which is the group of all permutations of the elements of the set $\{1,...,M\}$. For any permutation $\sigma\in\mathrm{Sym}(M)$ we define,
$S_{2}^{\sigma}(\tau):=\prod_{j=1}^{M}e^{\frac{\tau}{2}\hat{\mathcal{L}}_{\sigma(j)}}\prod_{k=M}^{1}e^{\frac{\tau}{2}\hat{\mathcal{L}}_{\sigma(k)}},$ (65)
which is a second order Trotter-Suzuki product formula whose exponentials are permuted by the permutation $\sigma\in\mathrm{Sym}(M)$. Using this, we can construct the approximation to the total channel $T_{\tau}$ by taking a convex combination of $S_{2}^{\sigma}$ for all $\sigma\in\mathrm{Sym}(M)$,
@@
where the $1/M!$ is present because $|\mathrm{Sym}(M)|=M!$. The following theorem shows the error bound and the bound on $N$ for the second order randomised Trotter-Suzuki product formula. The proof of this theorem is done in a similar way to the proof of the randomised product formulas in [11]. However, it also relies on the mixing lemma [30, 31], which as stated before is not applicable to open quantum systems, therefore we give a more direct proof.
@@
Given the generator $\mathcal{L}$ as in equation (10) of a quantum channel $T_{t}$ and some time $t\geq 0$. Define the second order randomised product formula as in equation (66). Let $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$ and choose $N$ such that $Mt\Lambda/N\leq 1$. Then,
$\left\|T_{t}-S_{2}^{(ran)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\frac{e(2\Lambda t)^{3}M^{2}}{N^{2}},$ (67)
where
$\epsilon\geq\frac{e(2\Lambda t)^{3}M^{2}}{N^{2}},$ (68)
and
$N\geq\frac{e^{1/2}(2\Lambda t)^{3/2}M}{\epsilon^{1/2}}.$ (69)
In proving the following theorem we will need to Taylor expand the formula $S_{2}^{(ran)}$. However this may be a complicated and challenging task to do directly. Instead, we want to consider what an arbitrary order term looks like in the expansion. The following lemma, which we shall call the randomisation lemma, will tell us what the $s$-th order non-degenerate term looks like in the expansion of $S_{2}^{(ran)}$, where $0\leq s\leq M.$ We use the word non-degenerate to describe a product of constituent generators $\hat{\mathcal{L}}_{k}$ which is pairwise different. To make out calculations easier we rewrite the second order randomised formula $S_{2}^{(ran)}$ in the following general way
@@
The last sum in equation (4) is a permutation of all pairwise different $m_{1},...,m_{s}$ and we observe that for a fixed $m_{1},...,m_{s}$ there are (Ms)!(M-s)! ways we can permute the rest of the indicies so that $m_{1},...,m_{s}$ is unchanged. Therefore, we remove this sum and add a factor (Ms)!(M-s)!, leading to the following expression for the $s$-th order non-degenerate term:
@@
To derive the bound on the norm of the s-th order degenerate term, we use a combinatorial argument. The complete set of $s$-th order terms can be partitioned into two disjoint sets: those with pairwise different indices (non-degenerate) and those with at least one repeated index (degenerate). The bound on the norm of the entire s-th order term is the sum of the norm bounds of all constituent terms. Our approach is to subtract the contribution of the non-degenerate terms from this total bound. The bound for the non-degenerate part is obtained by counting only the terms with distinct indices. The remaining contribution must therefore come from the degenerate terms. This counting-based approach is valid and is used in related literature on randomized simulation [11].
We first consider the degenerate term of the ideal evolution $T_{\tau}$. The upper bound on the norm of the degenerate term is given by,
$\left\|\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{degenerate}\end{subarray}}^{M}\hat{\mathcal{L}}_{m_{1}}...\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond}$ $\leq\frac{\tau^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{degenerate}\end{subarray}}^{M}\left\|\hat{\mathcal{L}}_{m_{1}}\right\|_{\diamond}...\left\|\hat{\mathcal{L}}_{m_{s}}\right\|_{\diamond}$ (88)
$\leq\frac{\tau^{s}\Lambda^{s}}{s!}\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{degenerate}\end{subarray}}^{M},$ (89)
$=\frac{\tau^{s}\Lambda^{s}}{s!}\left[\sum_{m_{1},...,m_{s}=1}^{M}-\sum_{\begin{subarray}{c}m_{1},...,m_{s}=1\\ \text{non-degenerate}\end{subarray}}^{M}\right],$ (90)
$=\frac{\tau^{s}\Lambda^{s}}{s!}\left[M^{s}-M(M-1)...(M-(s+1))\right]$ (91)
where we use the fact that the number of degenerate terms in the sum is equal to the total number of terms minus the number of non-degenerate terms. We prove in a similar way that the upper bound on the norm of the degenerate term of $S_{2}^{(\mathrm{ran})}$.
@@
Given a generator $\mathcal{L}$ as in equation (10) and the ideal evolution $T_{\tau}$ for some small time step $\tau\geq 0$ as well as the second order randomised formula $S_{2}^{(ran)}(\tau)$ defined in equation (66). Let the error superoperator be defined as $E(\tau)=T_{\tau}-S_{2}^{(ran)}(\tau)$ and let $E_{s}(\tau)$ be the $s$-th order term in the Taylor expansion of $E(\tau)$. Then, the diamond norm of $E_{s}(\tau)$ is bounded for each order $s$ as,
$\left\|E_{s}(\tau)\right\|_{\diamond}\leq\begin{cases}0,\text{ for }0\leq s\leq 2,\\ \\ \frac{(\Lambda\tau)^{s}M^{s-1}}{(s-2)!},\text{ for }s>2.\end{cases}$ (92)
@@
Now we need to obtain a bound for the factor $M^{s}-M(M-1)...(M-(s+1))]$. Using [11] we can obtain the following bound
@@ 5 The QDRIFT Channel For Simulating OQS
Given the generator $\mathcal{L}$ as in equation (8) of a quantum channel $T_{t}$ and some time $t\geq 0$. Let $\Omega:=\max_{k}\left\|\mathcal{L}_{k}\right\|_{\diamond}$ and choose $N$ such that $t\Gamma\Omega/N\leq 1$. Then,
$\left\|T_{t}-\left(\mathcal{E}^{(QD)}_{\omega}\right)^{N}\right\|_{\diamond}=\left\|T_{t}-\left(\sum_{k=1}^{M}p_{k}e^{\omega\mathcal{L}_{k}}\right)^{N}\right\|_{\diamond}\leq\frac{et^{2}\Gamma^{2}\Omega^{2}}{N}$ (109)
@@
$\epsilon\geq\frac{et^{2}\Gamma^{2}\Omega^{2}}{N}$ (110)
@@
$N\geq\frac{et^{2}\Gamma^{2}\Omega^{2}}{\epsilon}.$ (111)
@@ 6.1.1 Implementation of $S_{1}^{(ran)}$ with CS
To implement the first order randomised product formula $S_{1}^{(ran)}(\tau)^{N}$, we need $N$ applications of $S_{1}^{(ran)}(\tau)$. However, the definition of $S_{1}^{(ran)}(\tau)$ in (30) tells us that when we apply it to a state $\rho(0)$ it will apply the channel $S_{1}^{\rightarrow}(\tau)$ with probability $1/2$ and $S_{1}^{\leftarrow}(\tau)$ with probability $1/2$. This gives us a way to construct a gate set for $S_{1}^{(ran)}(\tau)$. If we define some random variable $j_{l}\in\{0,1\}$ for l=1,Nl=1,...N where $p(j_{l}=0)=1/2=p(j_{l}=1)$ then, by assigning $S_{1}^{\rightarrow}(\tau)\equiv S_{1}^{(0)}(\tau)$ and $S_{1}^{\leftarrow}(\tau)\equiv S_{1}^{(1)}(\tau)$, we can construct a set of operations, denoted by $G_{1}^{(ran)}$, by iteratively sampling each $j_{l}$ and appending $S_{1}^{(j_{l})}(\tau)$ to $G_{1}^{(ran)}$. The gate set can then be applied to an initial state $\rho(0)$ to output the state $\tilde{\rho}(t)$ which approximates $\rho(t)$ to an precision $\epsilon=(Mt\Lambda)^{3}/3N^{2}$. The pseudo code for the algorithm that can construct this gate set is shown in Algorithm 1.
@@
Input: A list of terms from the generator $\mathcal{L}=\sum_{k=1}^{M}\hat{\mathcal{L}}_{k}$ i.e. $\{\hat{\mathcal{L}}_{1},...,\hat{\mathcal{L}}_{M}\}$. A classical oracle function called SAMPLE() that returns a value $j_{l}\in\{0,1\}$ from the distribution $p(j_{l}=0)=1/2$ and $p(j_{l}=1)=1/2$. A target precision $\epsilon$ and a simulation time $t\geq 0$.
@@
[…]
@@
We want to obtain the gate complexity for the circuit implemented by $G_{1}^{(ran)}$. To do this, we count the number of simple channels, which in this case is the number of exponentials, $\exp(\tau\hat{\mathcal{L}}_{k})$, in the set of operations. Since there are $\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ TS product formulas in $G_{1}^{(ran)}$, each containing $M$ exponentials, the gate complexity scales as $O(M^{5/2}(t\Lambda)^{3/2}/\sqrt{3\epsilon})$.
@@ 6.1.2 Implementation of $S_{2}^{(ran)}$ with CS
To the implement the second order randomised formula $S_{2}^{(ran)}(\tau)^{N}$, we need $N$ applications of $S_{2}^{(ran)}(\tau)$ as in (4). By definition $S_{2}^{(ran)}(\tau)$ applies $S_{2}^{\sigma}(\tau)$ for some permutation $\sigma\in\mathrm{Sym}(M)$ with a probability $1/M!$. But since the sum in (4) is over all possible permutations in $\mathrm{Sym}(M)$ it is equally likely that we apply a $S_{2}^{\sigma}(\tau)$ for any permutation $\sigma$. So to construct a gate set that will implement $S_{2}^{(ran)}(\tau)^{N}$ we define $\sigma_{l}$, for l=1,2,,Nl=1,2,...,N, which will be some permutation from $\mathrm{Sym}(M)$. Then we define an oracle function called SAMPLE_PERMUTATION() which will sample from $\mathrm{Sym}(M)$, with every permutation in $\mathrm{Sym}(M)$ having a probability $1/M!$ of being sampled, and returning one of the permutations. We denote the gate set that we apply to initial state $\rho(0)$ as $G_{2}^{(ran)}$. Initially this set will be empty. Then, we iteratively obtain $\sigma_{l}$ using the function SAMPLE_PERMUTATION() i.e. $\sigma_{l}=$SAMPLE_PERMUTATION(), and append $S_{2}^{\sigma_{l}}$ to $G_{2}^{(ran)}$ repeating this process for l=1,2,,Nl=1,2,...,N. Then, we can apply $G_{2}^{(ran)}$ to some initial sate $\rho(0)$ and the output will be the state $\tilde{\rho}(t)$ which approximates $\rho(t)$ to a precision $\epsilon=(\Lambda t)^{3}M^{2}/N^{2}$. Algorithm 2, shows the pseudo code for how one can $G_{2}^{(ran)}$. Figure 3, shows the quantum circuit for appliyng each gate from $G_{2}^{(ran)}$ where each $\sigma_{l}$ is some permutation sampled from $\mathrm{Sym}(M)$. To see how the gate complexity for a circuit constructed with $G_{2}^{(ran)}$ scales, we count the number of exponentials. We will have $\lceil(\Lambda t)^{3/2}M/\sqrt{\epsilon}\rceil$ TS product formulas in $G_{2}^{(ran)}$ and for each of these there are $2M$ exponentials. The gate complexity will then scale as $O((\Lambda t)^{3/2}M^{2}/\sqrt{\epsilon})$.
@@
Input: A list of terms from the generator $\mathcal{L}=\sum_{k=1}^{M}\hat{\mathcal{L}}_{k}$ i.e. $\{\hat{\mathcal{L}}_{1},...,\hat{\mathcal{L}}_{M}\}$. A classical oracle function called SAMPLE_PERMUTATION() that returns a permutation from $\mathrm{Sym}(M)$, with each permutation having the probability $1/M!$ of being sampled. A target precision $\epsilon$ and a simulation time $t\geq 0$.
@@
[…]